OpenStudy (anonymous):

I need help on this question.. What is the simplified form of the quantity 1 over x minus 1 over y all over 1 over x plus 1 over y?

OpenStudy (anonymous):

PLEASE HELP ME I SWEAR THIS QUESTION IS IMPOSSIBLE

OpenStudy (anonymous):

@imqwerty @MikeyMaximum @AnimalLover8 @MikuxLuka @AnimalLover8 @arctic_blue808

OpenStudy (triciaal):

calm down

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ x }-\frac{ 1 }{ y } }{ \frac{ 1 }{ x}+\frac{ 1 }{ y } }\]

OpenStudy (triciaal):